OpenStudy (anonymous):

SOMEONE PLEASE HELP ME WITH THIS. A particle P moves in a straight line so that its displacement, S metres, from a fixed point O is given by S=4+15t-t³ where t is the time in seconds after passing through a point X on the line. --------------------------------------… Calculate the distance moved by the particle during the third second of its motion. Answer is 4.72m. Can any genius show me the steps?

ganeshie8 (ganeshie8):

familiar wid some calculus ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

not up to differential equations but yes i am familiar.

ganeshie8 (ganeshie8):

Good :) me too same. we can work it quickly just by using differentiation

ganeshie8 (ganeshie8):

first observe that the function has a zero slope between (2, 3) lets find this exact point, where the slope is zero

ganeshie8 (ganeshie8):

\( S=4+15t-t^3\) \(S' = 15 - 3t^2\) setting it to 0 and solving gives us \(t = \pm \sqrt{5}\)

ganeshie8 (ganeshie8):

\( \sqrt{5}\) is the time we're interested in

OpenStudy (anonymous):

beacuse at sqrt5 the particle is instantaneously at rest?

ganeshie8 (ganeshie8):

yup ! thats where, it changes its direction ! so for distance calculation, we need to calculate UP distance and DOWN distance both, between [2, 3]

ganeshie8 (ganeshie8):